# TreeLearn: A Comprehensive Deep Learning Method for Segmenting Individual Trees from Ground-Based LiDAR Forest Point Clouds

### v2

@@ 3 TreeLearn Pipeline
The goal of our TreeLearn pipeline is to identify the points belonging to trees in a 3D forest point cloud and then group these points into individual trees (Figure 2). More precisely, given a set of 3D points $P=\{p_{i}=(x_{i},y_{i},z_{i})\}_{i=1}^{N}$, our pipeline tackles the tasks of semantic and instance segmentation. The goal of semantic segmentation is to achieve a partition of all points into tree and non-tree points by assigning each point $p_{i}$ to one of the two classes. The non-tree class includes all points belonging to the forest ground and smaller vegetation such as bushes or trees smaller than 10 m. The tree class includes all points belonging to trees larger than 10 m. The goal of the instance segmentation step is to partition the tree points into $K$ (mutually exclusive) individual tree instances. The number of instances $K$ is not known a-priori and must be determined dynamically depending on the input data. The pipeline consists of the following five steps (Figure 2):
@@
    Use a neural network to predict two quantities for each point in a tile: (1) an offset vector pointing towards the x- and y-coordinates of the tree base (Figure 3a), and (2) a semantic score indicating the probability that a point belongs to a tree or not. This step is carried out separately for each tile generated in step 1.
@@
Running the complete TreeLearn pipeline on our 112 m by 103 m benchmark dataset takes about 180 minutes (3 GHz CPU; 16 physical cores of which one to two are used). The runtime is primarily determined by the CPU, since most time is spent in step one of the pipeline which includes feature calculation (30 minutes) and generating tiles and filtering them (100 minutes). GPU calculations require roughly 9 GB of memory and runtime is negligible. The runtime scales roughly linearly with the area of the plot.
3.1 Step 1: Generate Overlapping Forest Tiles
@@
The point clouds of an entire forest plot in our datasets contain around 20 000 0002000000020\,000\,00020 000 000 points. Processing a point cloud of this size at once using a neural network is infeasible due to memory restrictions. To address this issue, we employ a sliding window approach to process the forest. Predictions are generated for smaller, overlapping tiles that can be processed by the neural network. This approach is inspired by [52] who proposed tiling to process large input images. To create the tiles, we cut out rectangular sections with a fixed edge length of 30 m along the x- and y-axis from the original forest point cloud. The edge length along the z-axis is determined by the highest point. The tiles are generated in an overlapping way with a stride of 4 m. For each point within a tile, we calculate a feature that describes the local verticality of the point cloud [53], which is later used for clustering purposes (see Section 3.4).
Prior to generating the tiles, the lidar scans are subsampled with a voxel size of $10\text{cm}^{3}$, leaving only one point within a voxel. We argue that resolutions above 10 cm are not necessary for a large-scale task like tree segmentation. The tiles are further filtered using statistical and radius outlier removal [54]. The subsampling and filtering is designed to yield tiles with a reduced number of points to lower the computational demand, while at the same time preserving the complex structure of trees. Furthermore, the procedure is likely to produce similar results for data obtained by different laser scanners. For example, different laser scanner resolutions above 10 cm do not have an influence on the input data due to the subsampling procedure. This makes our method robust against changes in the input point cloud.
@@
Two things should be noted: first, due to the subsampling and filtering, our algorithm does not operate on the original point cloud. However, results on the original point cloud can be easily obtained by propagating the predictions from the filtered and subsampled point cloud to the original one, e.g. using k-nearest neighbors. Second, the tiling procedure can also be employed if the input point cloud is not rectangular. In fact, our method can be employed on arbitrarily shaped input point clouds.
3.2 Step 2: Predict Semantic Scores and Offsets
Figure 3: Visualizations regarding offset prediction. (a) depicts offsets for two example trees. The offset is the vector from a point towards the x- and y-coordinates of the tree base. The tree base is defined as the location of the trunk at a height of three meters. (b) visualizes the prediction area in relation to the whole tile. To make sure that the network has the necessary context information, predictions are only produced for the inner blue area of the tile.
@@ 3.2 Step 2: Predict Semantic Scores and Offsets
After having generated smaller processable point clouds, we use a neural network to predict two quantities for every point: first, we predict the probability that a point belongs to the semantic classes tree or non-tree. As these classes are mutually exclusive, this is a binary classification problem. Second, for every point we predict an offset vector [55] that points towards the x- and y-coordinates of the tree base. Specifically, the tree base is defined as the location of the trunk at a height of three meters. At this height, tree trunks are often further apart from neighboring close trunks than at the bottom. Figure 3a visualizes the offset vectors for a few example points.
Without any modifications, we use a well established neural network architecture that has been proven to be a powerful backbone for 3D scene understanding tasks [38, 39, 41]. The focus in this work is not on the network itself, but rather on its integration in a comprehensive segmentation pipeline. The network that we use for semantic and offset prediction takes as input a sparse voxel grid. First, the input space is partitioned into a 3D voxel grid where each voxel is $10~{}\text{cm}^{3}$ in size. Then, for each voxel it is checked whether it contains at least one point. If it does, the voxel is active and will be considered as part of the input to the network. Only active voxels are stored. All other voxels are ignored, hence the name ’sparse voxel grid’.
The network is a variant of the original U-Net introduced by [52]. The structure of a U-Net can be divided into two main steps. In the contraction step, features are computed in a hierarchical manner by gradually taking into account an increasingly larger context. In the expansion step, these features are propagated to the original points to arrive at rich voxel features that are used for semantic and offset prediction. For a more detailed explanation of the principles, we refer to the original paper [52]. Compared to the original U-Net, the network employed in this work follows the sparse convolution design proposed by [56]. Sparse convolutions differ from regular convolutions in that an activation in the output feature channel occurs only if the voxel at the center of the convolution kernel contains a point. This keeps the number of activations constant, which avoids excessive memory consumption. Furthermore, instead of just stacking convolutions, residual blocks are used to facilitate gradient flow [57]. The residual blocks consist of two 3 $\times$ 3 convolutions and a residual connection. Each contraction step of the U-Net consists of two residual blocks and a $2\times 2$ learned pooling layer implemented by a convolution with stride two. In this work, we employ seven contraction steps. During expansion, it is ensured that the sparse structure is preserved by only propagating activations to voxels that were active during the corresponding contraction step [58].
To predict the offset and semantic class, two separate heads are used, each of which take as input point features $F$. These point features $F=\{f_{1},\ldots,f_{N}\}\in\mathbb{R}^{N\times D}$ are obtained from the U-Net output by extracting for each of the $N$ original points the corresponding $D$-dimensional voxel feature. The semantic head maps onto semantic scores and consists of a multilayer perceptron (MLP) with two layers whose parameters are shared between points. Similarly, the offset head uses a shared three-layer MLP and maps onto offset predictions.
@@
Accurate offset prediction is possible only if the corresponding tree base is within the input tile, so the network can detect it. We therefore predict offsets only for the central 8 m $\times$ 8 m portion of each tile, while the network still receives the whole 30 m $\times$ 30 m tile as input (Figure 3b). This approach ensures that the network has sufficient context information around the points for which it makes predictions. Semantic predictions are also generated only for the inner square.
3.3 Step 3: Merge Tiles
@@
After semantic scores and offset vectors have been predicted for all tiles, the points from the inner squares along with their predictions are concatenated to obtain predictions for the full forest point clouds. No predictions are made for points outside the inner square. As the inner squares of neighboring tiles have an overlap of 4 m, multiple predictions for each point are generated. These predictions are then averaged, which reduces artifacts introduced by tiling and leads to smoother transitions between tiles. We calculate the projected coordinates $c_{i}=p_{i}+o_{i}$ by adding the predicted offsets ($o_{i}$) to the original points ($p_{i}$).
3.4 Step 4: Identify Tree Instances
@@
Since the goal is to arrive at tree instances, we only consider points with a higher probability to belong to the semantic class tree in this step. If the offset prediction for these points was perfect, the projected coordinates $c_{i}$ would have exactly the x- and y-coordinate of the corresponding tree base. In practice, the projected points that belong to the same tree still group together, but predictions are far from perfect. Projected points are often positioned in between neighboring high density areas that represent individual trees. These points form strings connecting the high density regions (Figure 2, see image at step 3). This is the result of prediction errors for points that cannot be clearly assigned to a tree, for example if the branches of two trees are heavily intertwined. In contrast, points belonging to the tree trunk have a much lower prediction uncertainty. Therefore, to obtain clearly separated tree clusters, we perform clustering only on points that fulfil the following two conditions: (1) the verticality feature of the point (see Section 3.1) has a sufficiently high value of $\tau_{\textrm{vert}}$, and (2) the absolute value of the z-component of the offset prediction is at most $\tau_{\textrm{off}}$. Condition (1) already separates trunk points from other tree points to a large degree. However, some points in tree crowns also have high local verticality. Therefore, condition (2) only selects those points that are sufficiently close to the tree base, thus ensuring that crown points are excluded from clustering.
@@
Individual tree clusters are identified by using the projected coordinates $c_{i}$ of the full forest point cloud that fulfil the verticality and offset condition above. These coordinates are used as input to a density-based clustering algorithm [59] following previous works [39, 38]. Specifically, an undirected graph is constructed where an edge between two points exists if their Euclidean distance is smaller than a predefined grouping radius $\tau_{\textrm{group}}$. Clusters are then obtained by identifying all connected components of the graph that consist of at least $\tau_{\textrm{min}}$ points (Figure 2, see image at step 4). The z-coordinate is ignored during clustering. The selection of the specific values for all hyperparameters $\tau$ used for instance identification is described in Section 4.3.
3.5 Step 5: Assign Remaining Points
@@
4 Training
The U-Net described in Section 3.2 was first trained using the noisy labels automatically obtained from Lidar360. The model obtained by training on noisy labels was then fine-tuned using the manually corrected clean labels.
4.1 Noisy Labels
@@
To generate the training data, we cropped out random areas of size 30 m $\times$ 30 m from the 18 forest plots that have only been automatically segmented. For these crops, subsampling and filtering was then performed according to the procedure described in Section 3.1. In total, 20 0002000020\,00020 000 crops were generated.
@@
To train the network, we calculate loss values on the predictions within the inner area of a crop, as displayed in Figure 3b. This way, the neural network almost always has enough context information. To supervise the offset prediction, we use the average L2-distance across tree points between the actual and predicted offset vector. For the semantic predictions, we use the binary cross entropy loss function and further scale it by a factor of 50 to obtain a similar magnitude to the offset loss. The offset and semantic loss are then summed to obtain the loss used for training. The U-Net was trained for 150 000150000150\,000150 000 iterations using the AdamW optimizer [60] with a weight decay of $10^{-3}$ and $\beta=[0.9,0.999]$. The batch size was set to 4. We further chose a cosine learning rate schedule [61] with a warm up period of $6250$ iterations and a maximum/minimum learning rate of $3\times 10^{-3}$/$5\times 10^{-6}$. Parameters pre-trained for 3D instance segmentation from [62] were used as initialization for the model. Network training was performed on a single Nvidia RTX A5000 and took roughly 40 hours.
4.2 Clean Labels

### v3

@@ 2.2 TreeLearn pipeline
The goal of our TreeLearn pipeline is to identify the points belonging to trees in a 3D forest point cloud and then group these points into individual trees. More precisely, given a set of 3D points $P=\{p_{i}=(x_{i},y_{i},z_{i})\}_{i=1}^{N}$, our pipeline tackles the tasks of semantic and instance segmentation. The goal of semantic segmentation is to achieve a partition of all points into tree and non-tree points by assigning each point $p_{i}$ to one of the two classes. The non-tree class includes all points belonging to the forest ground and smaller vegetation in the understory. The tree class includes all points belonging to trees. What exactly constitutes a tree for the model is determined by the annotations of the datasets used for training. The goal of the instance segmentation step is to partition the tree points into $K$ (mutually exclusive) individual tree instances. The number of instances $K$ is not known a-priori and must be determined dynamically depending on the input data. The pipeline consists of the following five steps (Figure 2):
@@
    Use a neural network to predict two quantities for each point in a tile: (1) an offset vector pointing towards the x- and y-coordinates of the tree base, and (2) a semantic score indicating the probability that a point belongs to a tree or not. This step is carried out separately for each tile generated in step 1.
@@
Running the complete TreeLearn pipeline on our 134 m by 125 m benchmark dataset (roughly 85 million points) took about 60 min (3 GHz CPU; 32 physical cores of which one to two are used). The runtime is primarily determined by CPU processing and can be roughly divided as following: subsampling and generating the tiles (28 min), averaging predictions (2 min), clustering (18 min), and propagation of the results to the original input point cloud (3 min). GPU calculations require roughly 9 GB of memory and runtime is comparably small (9 min). The runtime scales roughly linearly with the area of the plot. As research code, our implementation is not optimized for minimal runtime and resource usage.
2.2.1 Step 1: Generate overlapping forest tiles
@@
The point clouds of an entire forest plot in the datasets used contain tens of millions of points. Processing a point cloud of this size at once using a neural network is impractical due to memory restrictions. To address this issue, we employ a sliding window approach to process the forest. Predictions are generated for smaller, overlapping tiles that can be processed by the neural network. This approach is inspired by Ronneberger et al. [60] who proposed tiling to process large input images. To create the tiles, we cut out rectangular sections with a fixed edge length of 35 m along the x- and y-axis from the original forest point cloud. The edge length along the z-axis is determined by the highest point. The tiles are generated in an overlapping way with a stride of 4 m. For each point within a tile, we calculate a feature that describes the local verticality of the point cloud, which is used in later steps of the pipeline (see Section 2.2.2 and 2.2.4). It is calculated based on the relative magnitude of the z-component of locally computed Eigenvectors and ranges from a minimum of zero (no verticality) to a maximum of one (perfectly vertical). For a detailed description, we refer to Hackel et al. [61]. Prior to generating the tiles, the lidar scans are subsampled with a 10 cm voxel size, leaving only one point within a voxel. We argue that resolutions above 10 cm are not necessary for a large-scale task like tree segmentation. Subsampling also results in point clouds with the same resolution regardless of the original laser scanning resolution, provided it exceeds one point per 10 cm voxel.
@@
Two things should be noted: first, due to the subsampling, our algorithm does not operate on the original point cloud. However, results on the original point cloud are easily obtained by storing for each subsampled point the corresponding original points and propagating the segmentation labels at the end. This is implemented in our pipeline and has little effect on runtime. Second, the tiling procedure can also be employed if the input point cloud is not rectangular. In fact, our method can be employed on arbitrarily shaped input point clouds.
2.2.2 Step 2: Predict semantic scores and offsets
Figure 2: The pipeline for segmenting a forest point cloud. The circled numbers correspond to the five steps of the pipeline that are described in detail in Section 2.2. Apart from the model input and output, all images depict a top-view of the forest point cloud. The blue square on the image of step 1 shows the full tile, which is used as input to the network, while predictions are generated only for the inner red area. For illustrative purposes, a side-view of only the inner part is displayed as the model input, although the network receives the entire tile. The blue dots on the image of step 1 indicate that several tiles are being processed one after the other to obtain predictions for the full input point cloud. The colored points on the image of step 4 represent the initial clustering results which only take into account points that have a sufficiently high verticality and are close to the tree base (see Section 2.2.4 for details). The colored points on the image of step 5 represent the final predicted trees after postprocessing.
@@ 2.2.2 Step 2: Predict semantic scores and offsets
After having generated smaller processable point clouds, we use a neural network to predict two quantities for every point: first, we predict the probability that a point belongs to the semantic classes tree or non-tree. As these classes are mutually exclusive, this is a binary classification problem. Second, for every point we predict an offset vector [62] that points towards the x- and y-coordinates of the respective tree base. During training, we calculate the tree base coordinate as the mean of all points between 2.75 m and 3.25 m height that have a verticality feature value (see Section 2.2.1) of at least 0.6. This way, primarily trunk points are considered in the calculation of a tree base, while points from horizontal branches are ignored. Thus, the tree base can be intuitively understood as the location of the trunk at a height of three meters. The tree bases are then used to calculate the ground truth offset vectors, which are used as targets during training. During inference, the offset vectors are predicted since ground truth tree labels are not available. A height of 3 m for the tree base coordinate was chosen because, at this level, trunks are usually further apart from neighboring close trunks than at the bottom, which makes separation easier. This has already been leveraged in previous segmentation algorithms [13]. Figure 3a visualizes the offset vectors for a few example points.
@@
Accurate offset prediction is possible only if the corresponding tree base is within the input tile, so the network can detect it. For trees at the outer parts of the tile, there might only be a couple of branches available. It would be impossible to perform accurate offset prediction for these points. We therefore only predict offsets for the central 8 m $\times$ 8 m portion of each tile, while the network still receives the whole 35 m $\times$ 35 m tile as input (Figure 3b). Two things should be noted: (1) The predicted offset vector of a point within the central square can also point outside of the inner square in case that the trunk of the tree is located there. (2) Points within the central square of neighboring tiles will be projected to the same location if they belong to the same tree. This enables merging of the results (Section 2.2.3) and subsequent clustering of trees in a single step (Section 2.2.4). Semantic predictions are also generated only for the inner square.
The size of the central square (8 m $\times$ 8 m) in relation to the complete tile (35 m $\times$ 35 m) is based on a trade-off between tree base identifiability and computational efficiency: For points within the central square, the corresponding tree base is almost always located within the tile. Only one tree in all the forest point clouds used in this work is too wide to guarantee this for its outermost points. Further decreasing the size of the central square would ensure tree base identifiability in such extreme cases, but it would also reduce computational efficiency as more tiles would need to be processed. We therefore opted against this.
Without any modifications, we use a well established neural network architecture that has been proven to be a powerful backbone for 3D scene understanding tasks [43, 44, 46]. The focus in this work is not on the network itself, but rather on its integration in a comprehensive segmentation pipeline. The network that we use for semantic and offset prediction takes as input a sparse voxel grid. First, the input space is partitioned into a 3D voxel grid where each voxel has an edge length of 10 cm. Then, for each voxel it is checked whether it contains at least one point. If it does, the voxel is active and will be considered as part of the input to the network. Only active voxels are stored. All other voxels are ignored, hence the name ’sparse voxel grid’.
Figure 3: Visualizations regarding offset prediction. (a) depicts offsets for two example trees. The offset is the vector from a point towards the x- and y-coordinates of the tree base. The tree base is defined as the location of the trunk at a height of three meters. (b) visualizes the prediction area in relation to the whole tile. To make sure that the network has the necessary context information, predictions are only produced for the inner area of the tile. Offset prediction for points of the outer part of the tile is not possible in general since the corresponding tree bases might not be part of the tile.
The network is a variant of the original U-Net introduced by Ronneberger et al. [60]. The structure of a U-Net can be divided into two main steps. In the contraction step, features are computed in a hierarchical manner by gradually taking into account an increasingly larger context. In the expansion step, these features are propagated to the original points to arrive at rich voxel features that are used for semantic and offset prediction. For a more detailed explanation of the principles, we refer to the original paper [60]. Compared to the original U-Net, the network employed in this work follows the sparse convolution design proposed by Graham et al. [63]. Sparse convolutions differ from regular convolutions in that an activation in the output feature channel occurs only if the voxel at the center of the convolution kernel contains a point. This keeps the number of activations constant, which avoids excessive memory consumption. Furthermore, instead of just stacking convolutions, residual blocks are used to facilitate gradient flow [64]. The residual blocks consist of two 3 $\times$ 3 convolutions and a residual connection. Each contraction step of the U-Net consists of two residual blocks and a $2\times 2$ learned pooling layer implemented by a convolution with stride two. In this work, we employ seven contraction steps. During expansion, it is ensured that the sparse structure is preserved by only propagating activations to voxels that were active during the corresponding contraction step [65].
To predict the offset and semantic class, two separate heads are used, each of which take as input point features $F$. These point features $F=\{f_{1},\ldots,f_{N}\}\in\mathbb{R}^{N\times D}$ are obtained from the U-Net output by extracting for each of the $N$ original points the corresponding $D$-dimensional voxel feature. The semantic head maps onto semantic scores and consists of a multilayer perceptron (MLP) with two layers whose parameters are shared between points. Similarly, the offset head uses a shared three-layer MLP and maps onto offset predictions. The total number of trainable parameters of the network is roughly 30 million.
2.2.3 Step 3: Merge tiles
@@
After semantic scores and offset vectors have been predicted for the inner square of all tiles, these predictions and their corresponding points are concatenated to obtain predictions for the full forest point cloud. As the inner squares of neighboring tiles have an overlap of 4 m, multiple predictions for each point are generated. These predictions are then averaged, which reduces artifacts introduced by tiling and leads to smoother transitions between tiles. We calculate the projected coordinates $c_{i}=p_{i}+o_{i}$ by adding the predicted offsets ($o_{i}$) to the original points ($p_{i}$).
2.2.4 Step 4: Identify tree instances
@@
Since the goal is to arrive at tree instances, we only consider points with a higher probability to belong to the semantic class tree in this step. If the offset prediction for these points was perfect, the projected coordinates $c_{i}$ would have exactly the x- and y-coordinate of the corresponding tree base. In practice, the projected points that belong to the same tree still group together, but predictions are far from perfect. Projected points are often positioned in between neighboring high density areas that represent individual trees. These points form strings connecting the high density regions (Figure 2, see image at step 3). This is the result of prediction errors for points that cannot be clearly assigned to a tree, for example if the crowns of two trees are heavily intertwined. In contrast, points belonging to the tree trunk have a much lower prediction uncertainty. Therefore, to obtain clearly separated tree clusters, we perform clustering only on points that fulfil the following two conditions: (1) the verticality feature of the point (see Section 2.2.1) has a sufficiently high value of $\tau_{\textrm{vert}}$, and (2) the absolute value of the z-component of the offset prediction is at most $\tau_{\textrm{off}}$. Condition (1) already separates trunk points from other tree points to a large degree. However, some points in tree crowns also have high local verticality. Therefore, condition (2) only selects those points that are sufficiently close to the tree base, thus ensuring that crown points are excluded from clustering.
@@
Individual tree clusters are identified by using the projected coordinates $c_{i}$ of the full forest point cloud that fulfil the verticality and offset condition above. These coordinates are used as input to a density-based clustering algorithm [66] following previous works [44, 43]. Specifically, an undirected graph is constructed where an edge between two points exists if their Euclidean distance is smaller than a predefined grouping radius $\tau_{\textrm{group}}$. Clusters are then obtained by identifying all connected components of the graph that consist of at least $\tau_{\textrm{min}}$ points (Figure 2, see image at step 4). A separate visualization of the clustering results is also provided in Appendix B. The selection of the specific values for all hyperparameters $\tau$ used for instance identification is described in Section 2.4.
2.2.5 Step 5: Assign remaining points
@@
2.3 Training
The U-Net described in Section 2.2.2 was first trained using the noisy labels automatically obtained from Lidar360 (Section 2.3.1 below). The model obtained by training on noisy labels was then fine-tuned using the LAUTx, Wytham Woods and the For-Instance datasets (Section 2.3.2 below). Fine-tuning describes the process of using pre-trained model parameters as initialization when training with other datasets. This leads to a faster convergence compared to training from scratch. The performance of all trained models was benchmarked on the L1W and Wytham Woods datasets.
2.3.1 Pretraining with noisy labels
@@
To generate the training data, we cropped out random areas of size 35 m $\times$ 35 m from the 18 forest plots that have only been automatically segmented (Lidar360). These crops were then subsampled according to the procedure described in Section 2.2.1. In total, 20 0002000020\,00020 000 crops were generated.
@@
To train the network, we calculate loss values on the predictions within the inner area of a crop, as displayed in Figure 3b. This way, the neural network almost always has enough context information. To supervise the offset prediction, we use the average L2-distance across tree points between the actual and predicted offset vector. For the semantic predictions, we use the binary cross entropy loss function and further scale it by a factor of 50 to obtain a similar magnitude to the offset loss. The offset and semantic loss are then summed to obtain the loss used for training. The U-Net was trained for 150 000150000150\,000150 000 iterations using the AdamW optimizer [67] with a weight decay of $10^{-3}$ and $\beta=[0.9,0.999]$. The batch size was set to 2. We further chose a cosine learning rate schedule [68] with a warm-up period of $6250$ iterations and a maximum/minimum learning rate of $3\times 10^{-3}$/$5\times 10^{-5}$. Parameters pre-trained for 3D instance segmentation from Chen et al. [69] were used as initialization for the model. Network training was performed on a single Nvidia RTX A5000 and took roughly 40 hours.
2.3.2 Fine-tuning
The trained network can now already be used for segmentation. However, the performance is limited by the quality of the noisy training data. To make use of the smaller amount of manually labeled data, we perform fine-tuning of the pre-trained network with training crops generated from the LAUTx, FOR-Instance and Wytham Woods datasets.